OpenStudy (anonymous):

will fan and medal What is the value for ∆Soreaction for the following reaction, given the standard entropy values? 2 Al (s) + 3Cl2(g) 2AlCl3(s)

OpenStudy (anonymous):

-503 J/K +503 J/K -140 J/K +140 J/K

OpenStudy (anonymous):

plz help me

OpenStudy (photon336):

The sum of the entropy of the products minus the reactants: \[\Delta S = \sum (S_{products}-S_{reactants})\] \[2 Al (s) + 3Cl2(g) --> 2AlCl3(s)\] count number of moles of each reactants 2 moles Aluminum 3 moles of Chlorine The entropy for aluminum is 28 J/mol*k The entropy for chlorine gas is 223 J/mol*k products. 2 moles of aluminum tri chloride The entropy 111 j/mol*k Sum of the entropy for the reactants \[2*(28 \frac{ J }{ K mol }) + 3*(223\frac{ J }{ mol K }) = 725 \frac{ J }{ mol*K }\] products \[2*AlCl_{3}(\frac{ 111J }{ K mol }) = 222 \frac{ J }{ K mol } pro\] we subtract products - reactants. to get total entropy change \[222-725= -503 \frac{ J }{ k,mol }\] Because the entropy change is \[\Delta S < 0 \] this means that in terms of entropy, this reaction would not be favorable. but we would need to know the temperature as well as the sign of the enthalpy to really determine whether this reaction is spontaneous.

OpenStudy (photon336):

@ashleyZ